OpenStudy (firejay5):

Show work!!!!! 7. Navigation: The current in a river is 6 mph. In her motorboat Marissa can travel 12 miles upstream or 16 miles downstream in the same amount of time. What is the speed of her motorboat in still water?

OpenStudy (firejay5):

@ryan123345 Are you good at Math???

OpenStudy (kropot72):

Let s = speed in still water \[time=\frac{distance}{speed}\] \[\frac{12}{s-6}=\frac{16}{s+6}\] \[16s-96=12s+72\] Now just solve to find the value of s.

OpenStudy (firejay5):

s = 42 correct? @kropot72

OpenStudy (kropot72):

Yes, 42 mph is correct :)